# Exact Model Reduction for Continuous-Time Open Quantum Dynamics

### v2

@@ VI.2.2 Numerical model reduction and simulations
VI.3 Model reduction to study encoded dynamics
under non-ideal error models
@@ VI.3 Model reduction to study encoded dynamics
VI.3.1 Error dynamics on a three-bit subsystem code
As a concrete illustrative example, we consider the smallest infinite-distance noiseless subsystem code proposed in [62, 133] for protecting a qubit against arbitrary collective (permutation-invariant) noise using $N=3$ physical qubits. Let $s_{jk}\equiv\vec{\sigma}^{(j)}\cdot\vec{\sigma}^{(k)}=\sigma_{x}^{(j)}\sigma_{x}^{(k)}+\sigma_{y}^{(j)}\sigma_{y}^{(k)}+\sigma_{z}^{(j)}\sigma_{z}^{(k)}$, $j,k\in\{1,2,3\}$, denote the two-spin operators that are invariant under collective spin rotations, generated by collective angular momentum operators $\{J_{u},u=x,y,z\}$. One may show [133] that $P_{1/2}\equiv\frac{1}{2}\left[\mathbb{1}_{8}-\frac{1}{3}\left(s_{12}+s_{23}+s_{31}\right)\right]$ is the projector onto the 4-dimensional subspace $\mathcal{H}_{1/2}\subset\mathcal{H}=(\mathbb{C}^{2})^{\otimes 3}$ of states with total spin angular momentum eigenvalue $j=1/2$. If, as in Sec. VI.1, $\mathcal{S}_{3}$ denotes the permutation group on three objects, we have $s_{jk}\in\mathbb{C}\mathcal{S}_{3}$, while $J_{u}\in\mathbb{C}\mathcal{S}_{3}^{\prime}$. As one may verify, the following four observables in $\mathbb{C}\mathcal{S}_{3}$ then obey the algebraic properties of valid qubit observables [76]:
@@ VI.3.1 Error dynamics on a three-bit subsystem code
By implementing the algorithm described in Sec. IV.2, we may numerically determine the output algebra $\mathscr{O}$ that results from the combined dynamics in Eq. (60). Different possibilities arise, depending on the interplay between the nominal and error components (for illustration, see also Fig. 11, where we consider evolution starting from the initial state $\rho_{0}=\left|\psi_{0}\middle>\middle<\psi_{0}\right|$, with $\left|\psi_{0}\right>\equiv\left|0\right>_{L}\otimes\left|+1/2\right>_{Z}$ in Eq. (62)):
(a) $\Delta=1$, $\gamma_{j}=0$, $\forall j$
(b) $\Delta=10$, $\gamma_{j}=0$, $\forall j$
(c) $\Delta=2$, $\gamma_{1}=1$, $\gamma_{2}=\gamma_{3}=0$
(d) $\Delta=2$, $\gamma_{1}=\gamma_{2}=\gamma_{3}=1$
Figure 11: Evolution of the logical qubit corresponding to encoded observables $\{C_{q}\}$ [Eq. (61)], under error evolution due to permutation-symmetry-breaking and anisotropy in the Hamiltonian, plus local dephasing, $\gamma_{j}>0$, for the initial code state $\rho_{0}=(\left|0\middle>\middle<0\right|)_{L}\otimes(\left|+1/2\middle>\middle<+1/2\right|)_{Z}$ [Eq. (62)]. In (a), only permutation symmetry is broken, and the code remains unitarily noiseless under general collective noise. In (b), anisotropy is added with only collective dephasing, while in (c) and (d) local dephasing is also present. The relevant output algebra grows to the one of a qutrit in the last three cases. The solid lines indicate the trajectory of the reduced model, Eq.(60), while the empty dots denote the trajectory of the original one, Eq. (67). The model parameters are $\Gamma_{12}=\Gamma_{23}=1$, $\Gamma_{13}=0$, and $\kappa_{x}=\kappa_{y}=\kappa_{z}=1$ in (a), while $\kappa_{x}=\kappa_{y}=0$, $\kappa_{z}=1$ in (b), (c) and (d).
(i) $\Delta=1$, $\Gamma_{13}=0$, in the presence of general collective noise and no local dephasing ($\gamma_{j}=0,\forall j$, and $\kappa_{u}>0,\forall u$). In this case, $H$ no longer belongs to $\mathbb{C}\mathcal{S}_{3}^{\prime}$. However, since $H$ is still a linear combination of scalars under collective rotations, $\mathscr{E}_{\mathcal{O}}$ is preserved and information encoded in $\mathcal{O}$ evolves unitarily, The output algebra $\mathscr{O}=\mathscr{E}_{\mathcal{O}}$. Since unitary dynamics within the code space are reversible, $\mathcal{O}$ is unitarily noiseless [75, 68] [Fig. 11 (a)].
@@
(iv) $\Delta\neq 1$, $\Gamma_{13}=0$, in the presence of general collective noise and local dephasing ($\gamma_{j}>0$ for some $j$, and $\kappa_{u}>0,\forall u$). In this case, the output algebra becomes the full 16-dimensional algebra, $\mathscr{O}=\mathfrak{B}(\mathcal{H}_{\mathcal{C}})$. Interestingly, however, similar to the previous example in Sec. (VI.2), the observable space remains significantly smaller: here, we find $\dim(\mathscr{N}^{\perp})=16$. Thus, although no quantum MR is possible in this case, linear MR remains viable if desired.
VI.3.2 Reduced encoded dynamics
@@
${\check{O}}_{0}$ $\equiv\frac{1}{3}\left[\!\!\begin{array}[]{ccc}3&0&0\\ 0&1&\sqrt{2}\\ 0&\sqrt{2}&2\end{array}\!\right],\;{\check{O}}_{x}\equiv-\frac{1}{2}\left[\!\!\begin{array}[]{ccc}-1&1&\sqrt{2}\\ 1&{1}/{3}&{\sqrt{2}}/{3}\\ \sqrt{2}&{\sqrt{2}}/{3}&-{2}/{3}\end{array}\!\right],$
${\check{O}}_{y}$ $\equiv\frac{\sqrt{3}}{6}\left[\!\!\begin{array}[]{ccc}-3&-1&-\sqrt{2}\\ -1&1&\sqrt{2}\\ -\sqrt{2}&\sqrt{2}&2\end{array}\!\right],\;{\check{O}}_{z}\equiv\frac{i\sqrt{3}}{3}\left[\!\!\begin{array}[]{ccc}0&1&\sqrt{2}\\ -1&0&0\\ -\sqrt{2}&0&0\end{array}\!\right]\!\!.$
One can verify that these four observables $\{\check{O}_{q}\}$ obey su(2) commutation relationships and, in addition, $\check{O}_{q}^{2}=\check{O}_{0}$, for all $q$, as demanded for logical qubit observables [76].
@@
${\check{H}}=\mathcal{J}^{\dagger}(H)=$
$\left[\!\begin{array}[]{ccc}(-2-\Delta)\Gamma_{23}&(\Gamma_{13}-\Gamma_{12})\Delta&(\Gamma_{13}-\Gamma_{12})\sqrt{2}\\ (\Gamma_{13}-\Gamma_{12})\Delta&(2-\Delta)\Gamma_{23}&-(\Gamma_{13}+\Gamma_{12})\sqrt{2}\\ (\Gamma_{13}-\Gamma_{12})\sqrt{2}&-(\Gamma_{13}+\Gamma_{12})\sqrt{2}&(\Gamma_{12}-\Gamma_{13}-\Gamma_{23})\Delta\end{array}\!\right]\!.$
@@
${\check{L}}_{z,1}$ $=\gamma_{1}\left[\begin{array}[]{ccc}-1&0&0\\ 0&-1&0\\ 0&0&1\end{array}\right],$ ${\check{L}}_{z,2}$ $=\gamma_{2}\left[\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ 0&0&-1\end{array}\right],$
${\check{L}}_{z,3}$ $=\gamma_{3}\left[\begin{array}[]{ccc}0&-1&0\\ -1&0&0\\ 0&0&-1\end{array}\right],$ $\check{L}_{\text{id}}$ $=-\frac{\kappa_{z}}{2}\mathbb{1}_{3}.$
@@ VII Conclusions and outlook
(i) a linear model of provably minimal dimension, or
@@
Special emphasis in our analysis has been given to elucidate the role and usefulness of symmetries in enabling quantum MR. On the one hand, we have made precise the sense in which a strong symmetry of the Lindbladian results in a larger computational advantage than a weak one, with the improvement in system-size scaling being exponentially larger in principle (as we explicitly saw in a boundary-driven vs. an isolated or purely-dephasing XXZ chain). On the other hand and most importantly, the introduction of a new class of observable-dependent symmetries has allowed us to go beyond standard symmetry-based reductions, and show that exact MR is possible even in the absence of weak symmetries, as long as observable-dependent symmetries emerge for the observables of interest. It is natural to ask whether, in a dual manner, one could also introduce state-dependent symmetries, as unitary conjugations that only commute with the generator $\mathcal{L}$ when applied to certain initial conditions. This cannot be enough, however, to account for the reachable-based MR introduced in Sec. III, as even state-dependent symmetries would lead to projections onto standard unital algebras corresponding to their commutators. It would then remain to explain, from a symmetry-based viewpoint, how distorted algebras emerge. This is a first problem we leave for future work.
@@
Several other research questions are naturally prompted by the present investigation, some of which we highlight below.
@@
First, while we have considered the time-dependent expectations of the observable of interest as the quantities of interest for the simulation, our methods can be adapted to include the effect of the conditioning (state collapse) due to direct or indirect measurements, following the ideas presented for the discrete-time case in [70]. This would be especially interesting for obtaining reduced models to describe continuously monitored many-body quantum systems, which are being intensively investigated in the context of measurement-induced entanglement phase transitions [134, 135]. Likewise, since general time-local TP master equations have recently been shown to admit an unraveling in terms of a Markov process [136], this may also pave the way for extending MR to continuous-time systems described by genuinely non-Markovian, time-local master equations.
@@
Second, extensions to parametrized families of Lindbladian generators may be contemplated within our operator-algebraic framework, with appropriate modifications. Characterizing the level of MR robustly achievable for families of generators may enable a more complete understanding of Hilbert-space and operator-space fragmentation in open quantum systems, beyond the partial results available to date [94, 93]. Similarly, extensions to incorporate time-dependent control would be desirable, for instance to assess the effectiveness of methods such as dynamical decoupling in reducing decoherence. While combining Lindblad generators (e.g., adding an explicitly time-dependent Hamiltonian) is only possible under restrictive conditions [137], our approach would remain viable for structured environments, with Hamiltonian control applied on the system side at the level of the joint unitary dynamics, prior to MR. Such an extension to constructing exact reduced controlled models would complement, for instance, the (approximate) approach of [58], which allows for quantum control but, as noted, does not retain the quantum structure in general, or approaches based on cluster-expansion techniques as in [105], which are also approximate and purely numerical.
@@
Third, we have focused here on finite-dimensional systems, as they allow direct numerical implementation and validation of the proposed MR protocols using linear-algebraic tools. Nonetheless, the underlying ideas and methods would be well-worth extending to the infinite-dimensional setting, at the cost of added mathematical complexity [138]. This would open up applications to the large body of open quantum systems involving bosonic modes, which are of central relevance, for instance, to circuit quantum electrodynamics [139].
@@
Finally, while exact MR as we have considered is particularly appealing for both rigorous analysis and for obtaining easily interpretable reduced models, it may be too stringent for applications in many (if not most) realistic problems. The proposed framework can be adapted to include different kinds of approximate MRs. For instance, a first way of relaxing the exact setting would be to require that properties of interest are exactly reproduced only asymptotically in time, rather than along the full trajectory. This would suffice in many applications where only SS properties of the system are of interest, as in many quantum transport or SS criticality problems [19]. In moving to approximate MR, special care must be put in selecting an effective trade-off between accuracy and size of the target reduced model, while ensuring that it remains physically admissible. In this context, it would be especially interesting to determine whether modifications of the present methods may return a valid quantum model in settings where adiabatic elimination provably violates CP [140]. We look forward to report on some of these issues in future studies.
VIII Acknowledgments

### v3

@@ 6.2.2 Numerical model reduction and simulations
6.3 Model reduction to study encoded dynamics under non-ideal error models
@@
6.3.1 Error dynamics on a three-bit subsystem code
As a concrete illustrative example, we consider the smallest infinite-distance noiseless subsystem code proposed in [66, 135, 136] for protecting a qubit against arbitrary collective (permutation-invariant) noise using $N=3$ physical qubits. Let $s_{jk}\equiv\vec{\sigma}^{(j)}\cdot\vec{\sigma}^{(k)}=\sigma_{x}^{(j)}\sigma_{x}^{(k)}+\sigma_{y}^{(j)}\sigma_{y}^{(k)}+\sigma_{z}^{(j)}\sigma_{z}^{(k)}$, $j,k\in\{1,2,3\}$, denote the two-spin operators that are invariant under collective spin rotations, generated by collective angular momentum operators $\{J_{u},u=x,y,z\}$. One may show [136] that $P_{1/2}\equiv\frac{1}{2}\left[\mathbb{1}_{8}-\frac{1}{3}\left(s_{12}+s_{23}+s_{31}\right)\right]$ is the projector onto the 4-dimensional subspace $\mathcal{H}_{1/2}\subset\mathcal{H}=(\mathbb{C}^{2})^{\otimes 3}$ of states with total spin angular momentum eigenvalue $j=1/2$. If, as in Sec. 6.1, $\mathcal{S}_{3}$ denotes the permutation group on three objects, we have $s_{jk}\in\mathbb{C}\mathcal{S}_{3}$, while $J_{u}\in\mathbb{C}\mathcal{S}_{3}^{\prime}$. As one may verify, the following four observables in $\mathbb{C}\mathcal{S}_{3}$ then obey the algebraic properties of valid qubit observables [80]:
@@ 6.3.1 Error dynamics on a three-bit subsystem code
By implementing the algorithm described in Sec. 4.2, we may numerically determine the output algebra $\mathscr{O}$ that results from the combined dynamics in Eq. (60). Different possibilities arise, depending on the interplay between the nominal and error components (for illustration, see also Fig. 11, where we consider evolution starting from the initial state $\rho_{0}=\left|\psi_{0}\middle>\middle<\psi_{0}\right|$, with $\left|\psi_{0}\right>\equiv\left|0\right>_{L}\otimes\left|+1/2\right>_{Z}$ in Eq. (62)):
(i) $\Delta=1$, $\Gamma_{13}=0$, in the presence of general collective noise and no local dephasing ($\gamma_{j}=0,\forall j$, and $\kappa_{u}>0,\forall u$). In this case, $H$ no longer belongs to $\mathbb{C}\mathcal{S}_{3}^{\prime}$. However, since $H$ is still a linear combination of scalars under collective rotations, $\mathscr{E}_{\mathcal{O}}$ is preserved and information encoded in $\mathcal{O}$ evolves unitarily, with the output algebra $\mathscr{O}=\mathscr{E}_{\mathcal{O}}$. Since unitary dynamics within the code space are reversible, $\mathcal{O}$ is unitarily noiseless [79, 72] [Fig. 11 (a)].
@@
(iv) $\Delta\neq 1$, $\Gamma_{13}=0$, in the presence of general collective noise and local dephasing ($\gamma_{j}>0$ for some $j$, and $\kappa_{u}>0,\forall u$). In this case, the output algebra becomes the full 16-dimensional algebra, $\mathscr{O}=\mathfrak{B}(\mathcal{H}_{\mathcal{C}})$. Interestingly, however, similar to the previous example in Sec. (6.2), the observable space remains significantly smaller: here, we find $\dim(\mathscr{N}^{\perp})=16$. Thus, although no quantum MR is possible in this case, linear MR remains viable if desired.
6.3.2 Reduced encoded dynamics
@@
${\check{O}}_{0}$ $\equiv\frac{1}{3}\left[\!\!\begin{array}[]{ccc}3&0&0\\ 0&1&\sqrt{2}\\ 0&\sqrt{2}&2\end{array}\!\right],$
${\check{O}}_{x}$ $\equiv-\frac{1}{2}\left[\!\!\begin{array}[]{ccc}-1&1&\sqrt{2}\\ 1&{1}/{3}&{\sqrt{2}}/{3}\\ \sqrt{2}&{\sqrt{2}}/{3}&-{2}/{3}\end{array}\!\right],$
${\check{O}}_{y}$ $\equiv\frac{\sqrt{3}}{6}\left[\!\!\begin{array}[]{ccc}-3&-1&-\sqrt{2}\\ -1&1&\sqrt{2}\\ -\sqrt{2}&\sqrt{2}&2\end{array}\!\right],$
${\check{O}}_{z}$ $\equiv\frac{i\sqrt{3}}{3}\left[\!\!\begin{array}[]{ccc}0&1&\sqrt{2}\\ -1&0&0\\ -\sqrt{2}&0&0\end{array}\!\right]\!\!.$
These four observables $\{\check{O}_{q}\}$ obey su(2) commutation relationships and, in addition, $\check{O}_{q}^{2}=\check{O}_{0}$, for all $q$, as demanded for valid logical-qubit observables [80].
(a) $\Delta=1$, $\gamma_{j}=0$, $\forall j$
(b) $\Delta=10$, $\gamma_{j}=0$, $\forall j$
(c) $\Delta=2$, $\gamma_{1}=1$, $\gamma_{2}=\gamma_{3}=0$
(d) $\Delta=2$, $\gamma_{1}=\gamma_{2}=\gamma_{3}=1$
Figure 11: Time development of the logical qubit corresponding to encoded observables $\{C_{q}\}$ [Eq. (61)], under error evolution due to permutation-symmetry-breaking and anisotropy in the Hamiltonian, plus local dephasing, $\gamma_{j}>0$, for the initial code state $\rho_{0}=(\left|0\middle>\middle<0\right|)_{L}\otimes(\left|+1/2\middle>\middle<+1/2\right|)_{Z}$ [Eq. (62)]. In (a), only permutation symmetry is broken, and the code remains unitarily noiseless under general collective noise. In (b), anisotropy is added with only collective dephasing, while in (c) and (d) local dephasing is also present. The relevant output algebra grows to the one of a qutrit in the last three cases. The solid lines indicate the trajectory of the reduced model, Eq.(60), while the empty dots denote the trajectory of the original one, Eq. (67). The model parameters are $\Gamma_{12}=\Gamma_{23}=1$, $\Gamma_{13}=0$, and $\kappa_{x}=\kappa_{y}=\kappa_{z}=1$ in (a), while $\kappa_{x}=\kappa_{y}=0$, $\kappa_{z}=1$ in (b), (c) and (d).
@@
${\check{H}}=\mathcal{J}^{\dagger}(H)=\left[\!\begin{array}[]{ccc}(-2-\Delta)\Gamma_{23}&(\Gamma_{13}-\Gamma_{12})\Delta&(\Gamma_{13}-\Gamma_{12})\sqrt{2}\\ (\Gamma_{13}-\Gamma_{12})\Delta&(2-\Delta)\Gamma_{23}&-(\Gamma_{13}+\Gamma_{12})\sqrt{2}\\ (\Gamma_{13}-\Gamma_{12})\sqrt{2}&-(\Gamma_{13}+\Gamma_{12})\sqrt{2}&(\Gamma_{12}-\Gamma_{13}-\Gamma_{23})\Delta\end{array}\!\right]\!.$
@@
${\check{L}}_{z,1}$ $=\gamma_{1}\left[\begin{array}[]{ccc}-1&0&0\\ 0&-1&0\\ 0&0&1\end{array}\right],$
${\check{L}}_{z,2}$ $=\gamma_{2}\left[\begin{array}[]{ccc}0&1&0\\ 1&0&0\\ 0&0&-1\end{array}\right],$
${\check{L}}_{z,3}$ $=\gamma_{3}\left[\begin{array}[]{ccc}0&-1&0\\ -1&0&0\\ 0&0&-1\end{array}\right],$
$\check{L}_{\text{id}}$ $=-\frac{\kappa_{z}}{2}\mathbb{1}_{3}.$
@@ 7 Conclusions and outlook
(i) a linear model of provably minimal dimension,
or
@@
Special emphasis in our analysis has been given to elucidate the role and usefulness of symmetries in enabling quantum MR. On the one hand, we have made precise the sense in which a strong symmetry of the Lindbladian results in a larger computational advantage than a weak one, with the improvement in system-size scaling being exponentially larger in principle (as we explicitly saw in a boundary-driven vs. an isolated or purely-dephasing XXZ chain). On the other hand and most importantly, the introduction of a new class of observable-dependent symmetries has allowed us to go beyond standard symmetry-based reductions, and show that exact MR is possible even in the absence of weak symmetries, as long as observable-dependent symmetries emerge for the observables of interest. It is natural to ask whether, in a dual manner, one could also introduce state-dependent symmetries, as unitary conjugations that only commute with the generator $\mathcal{L}$ when applied to certain initial conditions. This cannot be enough, however, to account for the reachable-based MR introduced in Sec. 3, as even state-dependent symmetries would lead to projections onto standard unital algebras corresponding to their commutators. It would then remain to explain, from a symmetry-based viewpoint, how distorted algebras emerge. This is a first problem we leave for future work.
@@
Several other related research questions are naturally prompted by the present investigation, some of which we highlight below.
@@
(i) While we have considered the time-dependent expectations of the observable of interest as the quantities of interest for the simulation, our methods can be adapted to include the effect of the conditioning (state collapse) due to direct or indirect measurements, following the ideas presented for the discrete-time case in [74]. This would be especially interesting for obtaining reduced models to describe continuously monitored many-body quantum systems, which are being intensively investigated in the context of measurement-induced entanglement phase transitions [137, 138]. Likewise, since general time-local TP master equations have recently been shown to admit an unraveling in terms of a Markov process [139], this may also pave the way for extending MR to continuous-time systems described by genuinely non-Markovian, time-local master equations.
@@
(ii) Extensions to parametrized families of Lindbladian generators may be contemplated within our operator-algebraic framework, with appropriate modifications. Characterizing the level of MR robustly achievable for families of generators may enable a more complete understanding of Hilbert-space and operator-space fragmentation in open quantum systems, beyond the partial results available to date [96, 95]. Similarly, extensions to incorporate time-dependent control would be desirable, for instance to assess the effectiveness of methods such as dynamical decoupling in reducing decoherence. While combining Lindblad generators (e.g., adding an explicitly time-dependent Hamiltonian) is only possible under restrictive conditions [140], our approach would remain viable for structured environments, with Hamiltonian control applied on the system side at the level of the joint unitary dynamics, prior to MR. Such an extension to constructing exact reduced controlled models would complement, for instance, the (approximate) approach of [60], which allows for quantum control but, as noted, does not retain the quantum structure in general, or approaches based on cluster-expansion techniques as in [107], which are also approximate and purely numerical.
@@
(iii) We have focused here on finite-dimensional systems, as they allow direct numerical implementation and validation of the proposed MR protocols using linear-algebraic tools. Nonetheless, the underlying ideas and methods would be well-worth extending to the infinite-dimensional setting, at the cost of added mathematical complexity [141]. This would open up applications to the large body of open quantum systems involving bosonic modes, which are of central relevance, for instance, to circuit quantum electrodynamics [142].
(iv) While our reduced models are guaranteed to be in Lindblad form, they need not in general exhibit well-defined locality properties, as desirable on physical grounds. Adapting our framework so that locality constraints on observables or states (e.g., decay of correlations) are imposed in the MR procedure is a challenging problem well-worth investigating. In connection to the extension to state-dependent symmetries mentioned above, this may be possible to establish connections with operator-algebraic approaches to non-equilibrium many-body dynamics [143], whereby exact solutions for the behavior of arbitrary local quantities of interest at all times are shown to be determined by knowledge of a smaller set of special “pseudo-local” quantities. Reminiscent of our state-dependent symmetries, such operators need not obey the requisite symmetry requirements for all initial states.
@@
(v) While exact MR as we considered is particularly appealing for both rigorous analysis it may be too stringent for applications in many realistic problems. This calls for our proposed framework to be adapted to include different kinds of approximate MRs. In this venue, a first way of relaxing the exact setting would be to require that properties of interest are exactly reproduced only asymptotically in time, rather than along the full trajectory. This would suffice in applications where only SS properties of the system are of interest, as in many quantum transport or SS criticality problems [19] or where non-trivial, non-stationary dynamics emerges within the system’s center manifold in the long-time limit [144]. In moving to approximate MR, special care must be put in selecting an effective trade-off between accuracy and size of the target reduced model, while ensuring that it remains physically admissible. In this context, it would be interesting to determine whether modifications of the present methods may return a valid quantum model in settings where adiabatic elimination provably violates the CP property [145]. We look forward to report on some of these issues in future studies.
8 Acknowledgments